Soybean Oil is 16%protein; cornmeal is 8% protein. How many pounds of each should be mixed in order to get a 320 pound mixture that is 15% protein? How many pounds of cornmeal and how many pounds of soybean?
Answer by checkley77(12569)

.16X+.08(320-X)=.15*320
.16X+25.6-.08X=48
.08X=48-25.6
.08X=22.4
X=22.4/.08
X=280 POUNDS OF SOYBEANS IS USED.
320-280=40 POINDS OF CORNMEAL IS USED.
PROOF:
.16*280+.08*40=.15*320
44.8+3.2=48
48=48